# Toward Intelligent and Secure Cloud: Large Language Model Empowered Proactive Defense

### v1

@@ IV-A Experimental Setups
We simulate a prototype of LLM-PD with the support of GPT-4o mini, where five models are initially prompted to act as collector, analyzer, decision-maker, deployer, and feedback giver, respectively. Each model is constructed through a series of prompt engineering techniques, and Table I entails the input prompts for each component.
To evaluate our proposed architecture in defeating threats in cloud networks, we involve a DoS attack scenario. Specially, for SYN flooding and SlowHTTP attacks, we simulate an elastic service that can create up to 10 replicas, using a total pool of 100 pods. Each pod supports 256 connections and a maximum memory utilization of 100. The initial setup activates 5 replicas, each comprising 10 pods. Moreover, we also discuss Memory DoS, a co-resident attack that specifically targets cloud infrastructure. Here we model a cluster infrastructure with 5 racks, each containing 10 physical machines, and each machine capable of hosting up to 10 Virtual Machines (VMs). The maximum number of memory contention per VM is limited to 100 with runtime capped at 100 time steps. We conduct 10 episodes per test, in which an episode concludes when the system status remains stable for 5 consecutive steps.
In addition, we extensively compare our proposed method with several well-known existing schemes, including DQN [6], Actor-Critic (AC) [13], and Proximal Policy Optimization (PPO) [14]. For fairness in our experiments, we adopt the same configurations and environments for training each scheme with the discount factor set to 0.98. The learning rate is 0.001 for the Q-network in DQN and the policy network in AC and PPO, and 0.01 for the value network in AC and PPO. All networks are shallow, with a single hidden layer containing 256 neurons. All the experimental results are collected from 200 independent tests, which are performed to mitigate environmental uncertainties.
@@ IV-B Results
Fig. 2 presents the performance when defending against three attacks within different attack strengths, where the $x$-axis represents the number of episodes, the $y$-axis represents the average surviving rate under the current episode, and $n$ denotes the number of attackers. It is important to note that the average survival rate is measured by the probability of successfully mitigating an attack at each time step after a defensive action is taken to protect the system. Additionally, an episode refers to a complete operation cycle of LLM-PD architecture, and each episode retains the memory of feedback obtained from previous episodes. As we can see, the average surviving rate gradually increases for all attacks and strengths when running tests for more episodes, which implies the effectiveness of the generated policy, the correct deployment of defense mechanisms, and especially, the significant increase brought by the learned experience. Although more attackers pose a challenge and slightly decrease its performance, our solution still exhibits high surviving rates of 88.8%, 92.1%, and 93.5% when defending against attacks launched from 50 attackers due to the guidance from the LLM-based agents in our defense architecture.
@@
We also present the number of average steps needed for successful defense against these attacks within 50 attackers in Table II, where $e$ denotes the number of episodes. There is an observation that the steps for mitigating SYN Flooding, SlowHTTP, and Memory DoS attacks is 9.79, 9.55, and 16.3, respectively, when $e=2$. This implies the difficulty of defeating different attacks is significantly different when equipped with less knowledge and experience. As the episode progresses, the accumulation of knowledge allows our method to respond with fewer steps and reduced response times, reducing the number of steps needed to successfully defend against various attacks. Specifically, the number of steps required converges to 7.51, 7.23, and 7.53 when $e=10$. This can be attributed that LLM-PD avoids ineffective strategies and actions based on previous memory, thus promoting efficient resolution of diversified threats.
@@
In addition, we compare LLM-PD with three previously discussed solutions in terms of success rate, and all methods are trained or prompted with the SYN Flooding scenario of 10 attackers in this comparison. It is worth noting that the success rate is a relatively broad indicator, focusing solely on whether the attack is ultimately resisted successfully, without considering whether all intermediate processes are optimal. As illustrated in Fig. 3, all methods exhibit more than 97.4% of success rate, and especially, our solution achieves the highest performance when there are 10 attackers. Nevertheless, unlike existing approaches, LLM-PD encapsulates a vast spectrum of knowledge and a rich understanding from previous experience that enables it to act without additional or specific training. Not only that, when there are more attackers involved, LLM-PD exhibits remarkable adaptability, remaining high success rate while other approaches significantly reduce to 48.5%, 24.5%, and 31.5% when dealing with 50 attackers. This observation implies that LLM-PD can be easily and swiftly adapted to new tasks or scenarios with relatively minimal additional training, offering a significant advantage over other methods that may require extensive retraining.
@@
V Challenges and Future Directions
V-A Developing Explainable LLMs in the Security Domain
LLMs are notoriously complex black-box systems, where the intricate nature of their design poses significant challenges to interpretation and understanding. As these models become integrated into design, management, and control of the defense, and are entrusted with authoritative and decision-making roles, it is necessary to comprehend how they arrive at their conclusions and the rationale behind their decisions. In addition, understanding the logic behind can promote us to optimize and improve them, as well as trusting the behaviors of LLM agents. Therefore, developing explainable LLMs is essential for building user trust, ensuring a responsible and reasonable use of these models in the security domain.
V-B Training and Updating Fully Automatic LLM agents
The development of fully-automatic, security-oriented LLM agents necessitates the integration of diverse capabilities in multiple techniques to achieve effective and efficient end-to-end protection, promoting a paradigm shift from human-assisted operations to the full replacement of security experts. Moreover, a significant challenge lies in ensuring the model’s long-term relevance. The rapid evolution of network technologies and security tools demands continuous updates to the training data. In this context, maintaining and updating the model in a resource-efficient manner is a critical issue for future research and efforts.
V-C Designing LLM Empowered Secure Network Components
The proposed LLM-based proactive defense architecture is integrated into the existing cloud environment through additional attachment. However, the network components and defense mechanisms are predominantly implemented by humans, which may introduce inherent flaws and maintenance challenges. Consequently, a significant direction for future research will be the design of LLM empowered network components that are flexible to adapt to various network architectures and incorporate built-in security features. Among them, compatibility and reliability will be key issues, that is, the newly generated components need to be able to seamlessly integrate with existing infrastructure, and be able to operate stably under high concurrency and load scenarios.
@@ VI Conclusion
The development of LLM is promising for tackling challenges associated with mitigating cyberattacks. In this paper, we first introduce an innovative proactive defense architecture empowered by LLM to facilitate security improvement and threat mitigation within the cloud environment. By leveraging the advanced capabilities of LLM in data collection, security analysis, task inference, defense deployment, and effectiveness evaluation, our method is capable of thoroughly analyzing the security situation, efficiently executing suitable actions, and continuously evolving itself to adapt to varying and complex attack scenarios. Then a detailed case study on three kinds of DoS attacks is discussed. Experimental results demonstrate that the proposed architecture improves effectiveness and efficiency of defense than the state-of-the-art methods. Finally, we highlight future directions and open challenges to further enhance cloud security.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62202097 and No. 62072100, in part by China Postdoctoral Science Foundation under Grant No. 2024T170143 and No. 2022M710677, in part by Jiangsu Funding Program for Excellent Postdoctoral Talent under Grant No. 2022ZB137, and in part by the Southeast University-China Mobile Research Institute Joint Innovation Foundation under Grant No. CMYJY-202100163.

### v2

@@ IV-A Experimental Setups
It’s important to note that our LLM-PD architecture is model-agnostic and can be implemented with various LLMs. For our current implementation, we simulate a prototype of LLM-PD with GPT-4o mini due to its effective balance of reasoning capabilities and computational efficiency. Each agent is built through a series of prompt engineering techniques, and Table I presents the input prompts for each component.
To evaluate our proposed architecture in defeating threats, we implement a DoS attack scenario discussed in Section III-F. Specifically, for SYN flooding and SlowHTTP attacks, we simulate an elastic service that can create up to 10 replicas, using a total pool of 100 pods. Each pod supports 256 connections and a maximum memory utilization of 100. The initial setup activates 5 replicas, each comprising 10 pods. Additionally, we discuss Memory DoS, a co-resident attack that targets cloud infrastructure. Here we model a cluster with 5 racks, each containing 10 physical machines, and each machine capable of hosting up to 10 Virtual Machines (VMs). The maximum number of memory contention per VM is limited to 100 with runtime capped at 100 time steps. We conduct 10 episodes per test, in which an episode concludes when the system status remains stable for 5 consecutive steps.
We compare our proposed method with several well-known existing schemes, including DQN [6], Actor-Critic (AC) [13], and Proximal Policy Optimization (PPO) [14]. We adopt the same configurations and environments for training each scheme with the discount factor set to 0.98. The learning rate is 0.001 for the Q-network in DQN and the policy network in AC and PPO, and 0.01 for the value network in AC and PPO. All networks are shallow, with a single hidden layer containing 256 neurons. Since our method does not rely on training but rather on its own learning and evolution capabilities, these attacks are unseen by LLM-PD. To ensure statistical reliability and mitigate environmental uncertainties, we conduct 200 independent tests with initialization prior to each run.
@@ IV-B Results
Fig. 2 presents the performance when defending against three attacks within different attack strengths, where the $x$-axis represents the number of episodes, the $y$-axis represents the average surviving rate under the current episode, and $n$ denotes the number of attackers. Note that the average survival rate is measured by the probability of successfully mitigating an attack at each time step after a defensive action is taken to protect the system. Additionally, an episode refers to a complete operation cycle of LLM-PD, and each episode retains the memory of feedback obtained from previous episodes. The results show that the average surviving rate gradually increases for all attacks and strengths when running tests for more episodes, which implies the effectiveness of the generated policy, the correct deployment of defense mechanisms, and especially, the significant increase brought by the learned experience. Although more attackers pose a challenge and slightly decrease its performance, our solution still exhibits high surviving rates of 88.8%, 92.1%, and 93.5% when defending against attacks launched from 50 attackers due to the guidance from the LLM-based agents.
@@
We also present the number of average steps needed for successful defense against these attacks within 50 attackers in Table II, where $e$ denotes the number of episodes. There is an observation that the steps for mitigating SYN Flooding, SlowHTTP, and Memory DoS attacks are 9.79, 9.55, and 16.3, respectively, when $e=2$. This implies that the difficulty of defeating different attacks varies significantly when equipped with less knowledge and experience. As the episode progresses, the accumulation of knowledge allows our method to respond with fewer steps and reduced response times to successfully defend against various attacks. Specifically, the number of steps required converges to 7.51, 7.23, and 7.53 when $e=10$. This can be attributed to the fact that LLM-PD avoids ineffective strategies and actions based on previous memory, thus promoting efficient threat mitigation.
@@
In addition, we compare LLM-PD with three previously discussed solutions in terms of success rate, and all methods are trained or prompted with the SYN Flooding scenario of 10 attackers in this comparison. It is worth noting that the success rate is a relatively broad indicator, focusing solely on whether the attack is ultimately resisted successfully, without considering whether all intermediate processes are optimal. As illustrated in Fig. 3, all methods exhibit more than 97.4% success rate, and especially, our solution achieves the highest performance when there are 10 attackers. Nevertheless, unlike existing approaches, LLM-PD encapsulates a vast spectrum of knowledge and a rich understanding from previous experience that enables it to act without additional or specific training. Not only that, when there are more attackers involved, LLM-PD exhibits remarkable adaptability, maintaining a high success rate while other approaches significantly reduce to 48.5%, 24.5%, and 31.5% when dealing with 50 attackers. This observation implies that LLM-PD can easily and swiftly adapt to new tasks or scenarios with relatively minimal additional training, offering a significant advantage over other methods that may require extensive retraining.
V Challenges and Future Directions
V-A Developing Context Management for Long-term Memory
The current implementation of LLM-PD faces limitations in maintaining contextual awareness across extended security incidents. While the architecture includes a sliding window mechanism, LLMs inherently struggle with managing long-term dependencies and correlating events across extended timeframes, which is crucial for detecting sophisticated advanced persistent threats (APTs) that unfold over weeks or months. Future work should explore specialized context management architectures for security applications, enabling more effective correlation of events across extended time periods. Developing temporal reasoning capabilities specific to attack progression patterns would significantly improve detection and mitigation of multi-stage, long-running security threats.
@@ V-A Developing Context Management for Long-term Memory
V-B Designing Resource-Efficient LLM Based Security Solution
LLM-PD faces practical deployment challenges related to computational resource demands and integration with existing security infrastructure. High-capability LLMs require significant computational resources, which can introduce latency in time-sensitive scenarios and limit deployment in resource-constrained environments. Future efforts should focus on optimizing LLM-PD for resource efficiency through techniques like model distillation and specialized model architectures that maintain reasoning capabilities while reducing resource requirements. These advancements would transform LLM-PD from a promising research direction into a practical, widely deployable security solution for diverse cloud environments.
V-C Improving Robustness against Potential Adversarial Inputs
Another significant challenge specific to LLM-PD is its potential vulnerability to adversarial inputs designed to manipulate the defense system. Attackers familiar with LLMs might craft network traffic patterns or system behaviors specifically designed to trigger false interpretations or ineffective defense responses. To counter this, future research should focus on enabling LLM-PD to recognize and resist such manipulation attempts. This includes implementing robust verification mechanisms that cross-check LLM interpretations against multiple independent detection systems and developing techniques to identify inconsistencies between observed attack patterns and known attack taxonomies.
@@ VI Conclusion
The development of LLMs is promising for tackling challenges associated with mitigating cyberattacks. In this paper, we first introduce an innovative proactive defense architecture empowered by LLM to facilitate security improvement and threat mitigation within the cloud environment. By leveraging the advanced capabilities of LLM in data collection, security analysis, task inference, defense deployment, and effectiveness evaluation, our method is capable of thoroughly analyzing the security situation, efficiently executing suitable actions, and continuously evolving itself to adapt to varying and complex attack scenarios. Then, a detailed case study on three kinds of DoS attacks is discussed. Experimental results demonstrate that the proposed architecture improves effectiveness and efficiency of defense compared to the state-of-the-art methods. Finally, we highlight future directions and open challenges to further enhance cloud security.
@@ Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant No. 62202097, in part by China Postdoctoral Science Foundation under Grant No. 2024T170143 and No. 2022M710677, and in part by Jiangsu Funding Program for Excellent Postdoctoral Talent under Grant No. 2022ZB137.
@@
Yuyu Zhao [M’23] (yyzhao@seu.edu.cn) is a lecturer with the School of Cyber Science and Engineering at Southeast University.